OpenStudy (anonymous):

HELPP! Write the equation in slope-intercept form. What are the slope and y-intercept? -11x+9y=-12

OpenStudy (gw2011):

This should be in the form of: y=mx+b 9y=11x-12 y=(11/9)x-(12/9) y=(11/9)x-(4/3)

OpenStudy (anonymous):

Ahh. i think i got the answer. thanks.

OpenStudy (gw2011):

The slope = m = 11/9 y-intercept is equal when x = 0 y=(11/9)(0)-4/3 y=-(4/3)